OpenStudy (anonymous):

If f and g are inverse functions, then what is the composition of f and g equal to?

OpenStudy (aravindg):

Think of it like this: We are trying to find what fog gives us. g(x) gives you a result. f(g(x) Means f operates on g(x). f(x) being inverse of g undoes what g did. So you get back the input ie x.

OpenStudy (anonymous):

That makes me even more confused..

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

Oh then take an example. f(x)=x^2 f^-1=g(x)=root(x) f(g(z)=f(root(x))=(root x)^2=x Does that make sense?

OpenStudy (anonymous):

no, the ansswers either 1, 0, 2x, or x

OpenStudy (anonymous):

but i just dont know which one